Explain to me how to calculate average annual growth rate please?

In 1960Q3, real GDP per capita was $17,223. In 2017Q3, real GDP was per capita $52,605. What was the average annual growth rate over this period?

A 1.93%

B 3.6%

C 5.3%

D 1.97%
Answer #1

Solution:

Correct option is B. 3.6%.

From quarter 3 of 1960 to quarter 3 of 2017, there are 57 years. Over this time, the growth rate is:

Growth rate = (Final RGDP per capita - Initial RGDP per capita)*100/Initial RGDP per capita

Growth rate = ($52,605 - $17223)*100/17223

Growth rate = 35382*100/17223 = 205.43%

Average annual growth rate = Growth rate/number of years

Average annual growth rate = 205.43/57 = 3.604%

Thus, correct option is B) 3.6%